# Synergizing Foundation Models and Federated Learning: A Survey

### v1

@@
Abstract
The recent development of Foundation Models (FMs), represented by large language models, vision transformers, and multimodal models, has been making a significant impact on both academia and industry. Compared with small-scale models, FMs have a much stronger demand for high-volume data during the pre-training phase. Although general FMs can be pre-trained on data collected from open sources such as the Internet, domain-specific FMs need proprietary data, posing a practical challenge regarding the amount of data available due to privacy concerns. Federated Learning (FL) is a collaborative learning paradigm that breaks the barrier of data availability from different participants. Therefore, it provides a promising solution to customize and adapt FMs to a wide range of domain-specific tasks using distributed datasets whilst preserving privacy. This survey paper discusses the potentials and challenges of synergizing FL and FMs and summarizes core techniques, future directions, and applications. A periodically updated paper collection on FM-FL is available at https://github.com/lishenghui/awesome-fm-fl.
1 Introduction
@@
The landscape of Artificial Intelligence (AI) has been revolutionized by the emergence of Foundation Models (FMs) (Bommasani et al., 2021), such as BERT Devlin et al. (2019), GPT series Brown et al. (2020); OpenAI (2022, 2024), and LLaMA series Touvron et al. (2023a, b) in Natural Language Processing (NLP); ViTs Dosovitskiy et al. (2021) and SAM Kirillov et al. (2023) in Computer Vision (CV); CLIP Radford et al. (2021), DALL-E Ramesh et al. (2021), Gemini Google (2023), and GPT-4o in multimodal applications. These FMs have become pivotal in a myriad of AI applications across diverse domains. Their superb capability to generalize across tasks and domains stems from their pre-training on extensive datasets (Gunasekar et al., 2023), which imbues them with a profound understanding of language, vision, and multimodal data.
@@
While general-purpose FMs can leverage openly accessible data from the Internet, domain-specific FMs require proprietary data. It is, however, challenging to collect vast amounts of proprietary data and perform centralized pre-training or fine-tuning for domain-specific FMs, due to privacy restrictions Jo and Gebru (2020); GDPR (2016); CCPA (2023). Particularly in domains such as law, healthcare, and finance, where data is inherently privacy-sensitive, there is a pressing need for stringent privacy safeguards. Furthermore, given that data often constitutes a pivotal asset for enterprises, its widespread distribution is prohibitive. Consequently, there is an urgent need for novel strategies to handle data availability and facilitate model training, thereby unlocking the potential of domain-specific FMs whilst respecting data privacy.
@@
To address the challenges associated with data privacy in model training, Federated Learning (FL) (McMahan et al., 2017) has emerged as a promising paradigm. FL facilitates collaborative model training across decentralized clients without the need to share raw data, thus ensuring privacy preservation. Concretely, FL encompasses periodic interactions between the server and decentralized clients for the exchange of trainable model parameters without the requirement for private client data. Recognizing such a benefit, integrating FMs with FL presents a compelling solution for domain-specific FMs Zhuang et al. (2023); Yu et al. (2023d).
Despite the potential synergies between FL and FMs, the field is still nascent, lacking a comprehensive understanding of challenges, methodologies, and directions. This survey aims to bridge this gap by providing a thorough exploration of the integration of FMs and FL. We delve into the motivations and challenges of combining these two paradigms, highlight representative techniques, and discuss applications and future directions. By elucidating the intersection of FL and FMs, we aim to catalyze further research and innovation in this burgeoning area, ultimately advancing the development of privacy-aware, domain-specific FMs.
@@
The paper continues as follows: The next section introduces background on FMs and FL. Section 3 presents the motivation and challenges for synergizing FMs and FL. Section 4 highlights representative techniques. Section 5 explores the applications across various domains. Before concluding, we discuss representative future directions in Section 6.
2 Background
2.1 Foundation Models
@@
An FM is a model that can be adapted to a wide array of tasks through fine-tuning after initial pre-training Bommasani et al. (2021). The lifecycle of FMs typically involves pre-training on extensive generic data to establish the basis of their abilities Bubeck et al. (2023), followed by adaptation to downstream tasks such as domain-specific question answering Zhang et al. (2023e), and ultimately application in various domains.
@@
2.2 Federated Learning
FL McMahan et al. (2017) is a learning paradigm that enables a collection of clients to collaboratively learn a shared global model by leveraging their private datasets in a distributed manner, assisted by the coordination of a central server. The general goal of FL is to find a parameter set $\bm{\theta}$ that minimizes the following distributed optimization objective:
$\min\limits_{\bm{\theta}}F(\bm{\theta}):=\frac{1}{K}\sum_{k\in[K]}F_{k}(\bm{\theta}),$ (1)
@@ 2.2 Federated Learning
The most representative algorithms in the FL literature are the FedAvg-family algorithms McMahan et al. (2017); Reddi et al. (2021). The standard FedAvg involves periodic interactions between the server and decentralized clients to exchange trainable model parameters. In this process, each client independently trains the model on its local data and sends the model updates to a central server. The server aggregates these updates by computing their average to update the global model, which is subsequently redistributed to the clients for further iterations. Many variants have been proposed to tackle issues such as convergence and local data heterogeneity Diao et al. (2021). For example, FedProx Li et al. (2020) and FedDyn Acar et al. (2021) introduce regularizer terms to penalize client updates that are far away from the server model. A general framework FedOpt Reddi et al. (2021) unifies adaptive optimizers (Adam, Yogi, etc.) and demonstrates superior convergence speed when compared to the naive FedAvg.
FL offers an efficient privacy-preserving way to train models on large-scale and diverse data Kairouz et al. (2021), leading to its application across various domains such as healthcare Lincy and Kowshalya (2020); Rieke et al. (2020); Joshi et al. (2022), finance Chatterjee et al. (2023); Liu et al. (2023b), and smart cities Ramu et al. (2022); Pandya et al. (2023).
3 FM-FL: Motivation & Challenges
@@
3.1 Motivation
The integration of FMs and FL represents a compelling collaboration that leverages each other’s strengths to address their respective limitations, embodying a complementary relationship Zhuang et al. (2023); Li and Wang (2024).
@@ 3.1 Motivation
By leveraging data from a wide range of sources in a privacy-preserving manner, FL makes it possible to build models on sensitive data in specific domains, such as healthcare Lincy and Kowshalya (2020); Joshi et al. (2022); Rieke et al. (2020) and finance Chatterjee et al. (2023); Liu et al. (2023b). This enhances the diversity and volume of training data, improving model robustness and adaptability. Moreover, FL enables the integration of personal and task-specific data, allowing FMs to be customized for personal applications. For instance, Google has trained next-word-prediction language models on mobile keyboard input data with FL to improve user experience Xu et al. (2023); Bonawitz et al. (2021).
@@
By pre-training on large-scale generic data, FMs acquire essential knowledge and understanding capabilities Brown et al. (2020), providing multiple benefits to FL. Firstly, they benefit FL systems by offering advanced feature representations and learning capabilities from the outset. Secondly, leveraging the pre-learned knowledge of FMs can accelerate the FL process, enabling efficient and effective adaptation to specific tasks with minimal additional training. Thirdly, FMs’ powerful generative capabilities could help FL overcome the data heterogeneity challenge by synthesizing extra data, thus accelerating model convergence Huang et al. (2024).
3.2 Core Challenges
In this part, we discuss challenges emerging from the FM-FL marriage in three aspects: efficiency, adaptability, as well as trustworthiness.
@@ 3.2 Core Challenges
Efficiency challenges stem from the mismatch between the significant resource demands of FM training and the limited, heterogeneous system resources (e.g., mobile devices) within FL systems, such as communication bandwidth, computational power, and memory Su et al. (2023). The communication bottleneck of FL is induced by frequently exchanging training information between the server and clients over limited bandwidth channels Kairouz et al. (2021). The substantial number of parameters in FMs further exacerbates this burden, thus hindering the training process.
@@
Adaptability challenges arise from the adaptation of an FM to a specific downstream task (e.g., by fine-tuning) in FL settings. Key challenges include data heterogeneity and resource heterogeneity. Performance degradation in FL, attributed to heterogeneous data distributions among clients, is a well-recognized issue Kairouz et al. (2021); Li et al. (2022). A recent study Babakniya et al. (2023a) has shown that such performance penalty is even more substantial when fine-tuning FMs. For NLP tasks, data heterogeneity can manifest as variations in language, style, topic, or sentiment across datasets held by different clients. In multi-modal scenarios, the challenge is even more pronounced due to the inherent diversity in data types (e.g., text, images, and audio) Yu et al. (2023a). Addressing data heterogeneity involves not just identifying and measuring it but also developing algorithms that are robust to such diversity, ensuring that the model can learn effectively from varied data contributions without compromising on performance. In terms of resource heterogeneity, the memory and computational resources of the devices for different participants may be diverse Diao et al. (2021), which could cause delays for model synchronization and inactivation of some participants, i.e., stragglers, making it challenging to leverage the full potential of FMs in FL settings.
@@
Trustworthiness challenges emphasize the concerns regarding privacy, security, and ethical considerations in the lifecycle of FM-FL, from the pre-training and model adaptation to the application stages. We present two representative challenges from this perspective: (1) intellectual property: Intellectual Property (IP) protection in FM-FL primarily involves attributing ownership rights for both models and data. From the server’s perspective, broadcasting a pre-trained model to multiple nodes for fine-tuning poses IP protection and security risks (e.g., model theft), necessitating measures to safeguard IP rights and ensure model integrity Kang et al. (2024); (2) privacy leakage: Although FL does not immediately share data, studies have shown that it may not always guarantee sufficient privacy preservation Geiping et al. (2020), as model parameters (e.g., weights or gradients) may leak sensitive information to malicious adversaries Zhu et al. (2019). (3)Poisoning Attacks: FL systems are inherently vulnerable to attacks due to their wide attack surface and reliance on network communication Li et al. (2023b). Poisoning attacks are carried out by malicious participants, aiming to bias the global model to the desire of attackers.
@@
[…]
4.1), l1node, [Parameter-Efficient
Fine-Tuning, l2node, [Selective, l3node, [  RaFFM Yu et al. (2023c), FedBF Zhang et al. (2023f) , leaf] ] [Additive, l3node, [  FedCLIP Lu et al. (2023a), FedDAT Chen et al. (2024) , leaf] ] [Reparameterization-based, l3node, [  HetLoRA Cho et al. (2024), FedDPA Yang et al. (2024b), leaf ] ] ] [Model Compression, l2node, [Sparsification, l3node, [ PruneFL Jiang et al. (2023c), FLASH Babakniya et al. (2023b),leaf] ] [Quantization, l3node, [ FedSplitBERT Lit et al. (2022),leaf] ] ] [Zeroth-Order
Optimization, l2node, [  BAFFLE Feng et al. (2023b), FedZeN Maritan et al. (2023), FedKSeed Qin et al. (2024),
FwdLLM Xu et al. (2024a), ZooPFL Lu et al. (2023b), FedMeZO Ling et al. (2024),wide leaf] ] ] [Adaptability
4.2), l1node, [Domain-Centric, l2node, [Domain-Adaptive Pre-Training, l3node, [  FMTDA Yao et al. (2022), FEDBFPT Wang et al. (2023) , leaf ] ] [Multi-Domain Adaptation, l3node, [  FedAPT Su et al. (2024), DiPrompT Bai et al. (2024b), leaf ] ] ] [Client-Centric, l2node, [Personalization, l3node, [  FedDAT Chen et al. (2024), Fed-MNMT Liu et al. (2023d),leaf ] ] [Client Clustering, l3node, [  FedLFC Guo et al. (2024b), FL-TAC Ping et al. (2024),leaf ] ] ] [System-Centric, l2node, [Resource-Heterogeneous, l3node, [  FedRA Su et al. (2023), HetLoRA Cho et al. (2024),leaf ] ] [ Split Learning, l3node, [ FedBERT Tian et al. (2022), FedSplitX Shin et al. (2023b),leaf] ] ] ] [Trustworthiness
4.3), l1node, [IP Protection, l2node, [Watermarking, l3node, [  WAFFLE Tekgul et al. (2021), DUW Yu et al. (2023b),leaf ] ] [Black-Box Tuning, l3node, [  Fed-BBPT Lin et al. (2023), pFedGPT Rui et al. (2024),leaf ] ] ] [Privacy Protection, l2node, [Privacy-Preserving Techniques, l3node, [  DP-FTRL Xu et al. (2023), DP-LoRA Liu et al. (2023c) , leaf ] ] [Privacy Attack, l3node, [  FILM Gupta et al. (2022), DRA Zhang et al. (2024c) , leaf ] ] ] [Attack Robustness, l2node, [Poisoning Attacks, l3node, [ Fed-EBD Li et al. (2024c), leaf ] ] [Defense Techniques, l3node, [ ClippedClustering Li et al. (2023b), Fed-FA Zhang et al. (2023d), leaf ] ] ] ] ]
@@
Figure 1: Taxonomy of research in foundation models with federated learning.
4 Techniques
Recent work has begun to address challenges associated with adapting pre-trained FMs to specific downstream tasks in FL settings. In this section, we survey FM-FL techniques on three aspects, namely efficiency (Section 4.1), adaptability (Section 4.2), and trustworthiness (Section 4.3). As illustrated in Figure 1, we further refine them according to the key features of different methods.
4.1 Efficiency
There has been a considerable focus on developing resource-efficient approaches. This part describes techniques that improve resource efficiency.
4.1.1 Parameter-Efficient Fine-Tuning

### v2

@@
Abstract.
Over the past few years, the landscape of Artificial Intelligence (AI) has been reshaped by the emergence of Foundation Models (FMs). Pre-trained on massive datasets, these models exhibit exceptional performance across diverse downstream tasks through adaptation techniques like fine-tuning and prompt learning. More recently, the synergy of FMs and Federated Learning (FL) has emerged as a promising paradigm, often termed Federated Foundation Models (FedFM), allowing for collaborative model adaptation while preserving data privacy. This survey paper provides a systematic review of the current state of the art in FedFM, offering insights and guidance into the evolving landscape. Specifically, we present a comprehensive multi-tiered taxonomy based on three major dimensions, namely efficiency, adaptability, and trustworthiness. To facilitate practical implementation and experimental research, we undertake a thorough review of existing libraries and benchmarks. Furthermore, we discuss the diverse real-world applications of this paradigm across multiple domains. Finally, we outline promising research directions to foster future advancements in FedFM. Overall, this survey serves as a resource for researchers and practitioners, offering a thorough understanding of FedFM’s role in revolutionizing privacy-preserving AI and pointing toward future innovations in this promising area.
copyright: noneccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
A significant paradigm shift in Artificial Intelligence (AI) has been catalyzed by the rapid advancement of Foundation Models (FMs) (Bommasani et al., 2021), such as BERT (Devlin et al., 2019), GPT series (Brown and others, 2020; OpenAI, 2022, 2024), and LLaMA series (Touvron et al., 2023a, b) in Natural Language Processing (NLP); ViTs (Dosovitskiy et al., 2021) and SAM (Kirillov et al., 2023) in Computer Vision (CV); CLIP (Radford et al., 2021), DALL-E (Ramesh et al., 2021), Gemini (Google, 2023), and GPT-4o in multimodal applications. These FMs have become pivotal in a myriad of AI applications across diverse domains. The remarkable generalization capabilities of FMs are attributed to large-scale pre-training (Gunasekar et al., 2023), which enables them to learn robust representations of language, vision, and multimodal data. To leverage these representations for specific applications, adaptation techniques such as fine-tuning are widely adopted to tailor FMs to diverse downstream tasks (Zhao et al., 2023).
@@
While general-purpose FMs can leverage openly accessible data from the Internet, domain-specific FMs require proprietary data. It is, however, challenging to collect vast amounts of proprietary data and perform centralized pre-training or fine-tuning for domain-specific models, due to privacy restrictions (Jo and Gebru, 2020), i.e., General Data Protection Regulation (GDPR) (GDPR, 2016) and California Consumer Privacy Act (CCPA) (CCPA, 2023). Particularly in domains such as law, healthcare, and finance, where data is inherently privacy-sensitive, there is a pressing need for stringent privacy safeguards. Furthermore, given that data often constitutes a pivotal asset for enterprises, its widespread distribution is prohibitive. Consequently, there is an urgent need for novel strategies to handle data availability and facilitate model training, thereby unlocking the potential of domain-specific applications of FMs whilst respecting data privacy.
@@
To address the challenges of data privacy in model training, Federated Learning (FL) (McMahan et al., 2017) has emerged as a promising paradigm. FL facilitates collaborative model training across decentralized clients without sharing raw data, thereby preserving privacy. Specifically, FL involves periodic interactions between the server and decentralized clients to exchange trainable model parameters without requiring private client data. Recognizing such a benefit, integrating FMs with FL, often termed Federated Foundation Models (FedFM), presents a compelling solution for domain-specific FMs (Zhuang et al., 2023; Yu et al., 2024).
Table 1. Summary of previous surveys related to FedFM (: Included, : Partially included, : Not included)
Survey Year Taxonomy Technical Focus Library & Benchmark Application Future Direction
Efficiency Adaptability Trustworthiness
Zhuang et al. (2023) (2023)
Chen et al. (2023) (2023)
Yu et al. (2024) (2024)
Woisetschläger et al. (2024) (2024)
Hu et al. (2024) (2024)
Li and Wang (2024) (2024)
Yao et al. (2024b) (2024b)
Kang et al. (2025) (2025)
Wu et al. (2025b) (2025b)
Fan et al. (2025c) (2025c)
Ren et al. (2025) (2025)
Ours 2026
Driven by the surging interest in the intersection of FMs and FL, several surveys have investigated various dimensions of the FedFM paradigm. Table 1 compares our work with prior surveys in terms of coverage of key aspects. Most existing surveys focus on presenting the vision and challenges, but few analyze the overall system performance from multiple perspectives. Specifically, early contributions primarily articulated the motivation and outlined the key challenges of integrating FMs with FL (Zhuang et al., 2023; Yu et al., 2024; Chen et al., 2023). Subsequent work has considered specific topics but mostly focuses on efficiency (Woisetschläger et al., 2024; Hu et al., 2024; Fan et al., 2025c; Yao et al., 2024b), with very few studies on adaptability and trustworthiness (Kang et al., 2025; Wu et al., 2025b; Li and Wang, 2024). Moreover, these works rarely provide a unified taxonomy that captures the full landscape. Also, none of them considers all three important metrics on efficiency, adaptability, and trustworthiness holistically. To the best of our knowledge, the work by Ren et al. (2025) represents the only prior attempt to provide a comparatively comprehensive taxonomy across various dimensions of FedFM. However, this work does not consider adaptability a critical dimension; instead, it structures its taxonomy around the conventional FL lifecycle. We argue that this emerging paradigm necessitates a novel perspective to fit new scenarios and demands of FMs.
Recognizing gaps in the existing literature, we conduct a systematic review of FedFM research to provide a thorough exploration of system performance in the integration of FMs and FL. More specifically, we provide a comprehensive study and analysis of the existing work across three major performance perspectives: efficiency, adaptability, and trustworthiness. Our main contributions are detailed as follows:
  • We present a novel multi-tiered taxonomy for FedFM. This framework systematically classifies existing techniques based on their characteristics and modeling approaches. It serves as a foundation for FedFM methodologies, guiding the design of future solutions.
  • We discuss publicly available libraries and benchmarks, including their attributes, to make it easy for readers to get started with their research.
  • We introduce typical applications of FedFM in critical domains like NLP, Recommendation, and Healthcare.
@@
The remainder of this survey is organized as follows: Section 2 introduces background on FMs and FL. Section 3 presents the motivation and challenges for synergizing FMs and FL. Section 4, 5, and 6 review and summarize the recent progress from the three aspects of efficiency, adaptation, and trustworthiness. Section 8 explores the applications across various domains. Before concluding, we discuss representative future directions in Section 9.
2. Background
This section briefly describes the background of FMs and FL, respectively.
2.1. Foundation Models
@@
An FM is a model that can be adapted to a wide array of tasks through fine-tuning after initial pre-training (Bommasani et al., 2021). The lifecycle of FMs typically involves pre-training on extensive generic data to establish the basis of their abilities (Bubeck et al., 2023), followed by adaptation to downstream tasks such as domain-specific question answering (Zhang et al., 2023d), and ultimately application in various domains.
@@
2.2. Federated Learning
FL (McMahan et al., 2017) is a learning paradigm that enables a collection of clients to collaboratively learn a shared global model by leveraging their private datasets in a distributed manner, assisted by the coordination of a central server. The general goal of FL is to find a parameter set $\bm{\theta}$ that minimizes the following distributed optimization objective:
(1) $\min\limits_{\bm{\theta}}F(\bm{\theta}):=\frac{1}{K}\sum_{k\in[K]}F_{k}(\bm{\theta}),$
@@ 2.2. Federated Learning
The most representative algorithms in the FL literature are the FedAvg-family algorithms (McMahan et al., 2017; Reddi et al., 2021). The standard FedAvg involves periodic interactions between the server and decentralized clients to exchange trainable model parameters. In this process, each client independently trains the model on its local data and sends the model updates to a central server. The server aggregates these updates by computing their average to update the global model, which is subsequently redistributed to the clients for further iterations. Many variants have been proposed to tackle issues such as convergence and local data heterogeneity (Diao et al., 2021). For example, FedProx (Li et al., 2020) and FedDyn (Acar et al., 2021) introduce regularizer terms to penalize client updates that are far away from the server model. A general framework FedOpt (Reddi et al., 2021) unifies adaptive optimizers (Adam, Yogi, etc.) and demonstrates superior convergence speed when compared to the naive FedAvg.
FL offers an efficient privacy-preserving way to train models on large-scale and diverse data (Kairouz and others, 2021), leading to its application across various domains such as healthcare (Lincy and Kowshalya, 2020; Rieke et al., 2020; Joshi et al., 2022), finance (Chatterjee et al., 2023; Liu et al., 2023b), and smart cities (Ramu et al., 2022; Pandya et al., 2023).
3. FedFM: Motivation & Challenges
@@
3.1. Motivation
The integration of FMs and FL represents a compelling collaboration that leverages each other’s strengths to address their respective limitations, embodying a complementary relationship (Zhuang et al., 2023; Li and Wang, 2024).
@@ 3.1. Motivation
By leveraging data from a wide range of sources in a privacy-preserving manner, FL makes it possible to build models on sensitive data in specific domains, such as healthcare (Lincy and Kowshalya, 2020; Joshi et al., 2022; Rieke et al., 2020) and finance (Chatterjee et al., 2023; Liu et al., 2023b). This enhances the diversity and volume of training data, improving model robustness and adaptability. Moreover, FL enables the integration of personal and task-specific data, allowing FMs to be customized for personal applications. For instance, Google has trained next-word-prediction language models on mobile keyboard input data with FL to improve user experience (Xu et al., 2023; Bonawitz et al., 2021).
@@
By pre-training on large-scale generic data, FMs acquire essential knowledge and understanding capabilities (Brown and others, 2020), providing multiple benefits to FL. Firstly, they benefit FL systems by offering advanced feature representations and learning capabilities from the outset. Secondly, leveraging the pre-learned knowledge of FMs can accelerate the FL process, enabling efficient and effective adaptation to specific tasks with minimal additional training. Thirdly, FMs’ powerful generative capabilities could help FL overcome the data heterogeneity challenge by synthesizing extra data, thus accelerating model convergence (Huang et al., 2024b).
3.2. Core Challenges
In this section, we discuss challenges arising from FedFM across three major performance perspectives: efficiency, adaptability, and trustworthiness. They are considered important metrics for training and running the FMs across diverse downstream applications. We will explain them one by one below.
@@ 3.2. Core Challenges
Efficiency challenges stem from the mismatch between the significant resource demands of FM training and the limited, heterogeneous system resources (e.g., mobile devices) within FL systems, such as communication bandwidth, computational power, and memory (Su et al., 2024a). The communication bottleneck of FL is induced by frequently exchanging training information between the server and clients over limited bandwidth channels (Kairouz and others, 2021). The substantial number of parameters in FMs further exacerbates this burden, thus hindering the training process.
@@
Adaptability challenges arise from the adaptation of an FM to a specific downstream task (e.g., by fine-tuning) in FL settings. Key challenges include data heterogeneity and resource heterogeneity. Performance degradation in FL, attributed to heterogeneous data distributions among clients, is a well-recognized issue (Kairouz and others, 2021; Li et al., 2022). A recent study (Babakniya et al., 2023) has shown that such a performance penalty is even more substantial when fine-tuning FMs. For NLP tasks, data heterogeneity can manifest as variations in language, style, topic, or sentiment across datasets held by different clients. In multi-modal scenarios, the challenge is even more pronounced due to the inherent diversity in data types (e.g., text, images, and audio) (Yu et al., 2023a). Addressing data heterogeneity involves not just identifying and measuring it but also developing algorithms that are robust to such diversity, ensuring that the model can learn effectively from varied data contributions without compromising on performance. In terms of resource heterogeneity, the memory and computational resources of the devices for different participants may be diverse (Diao et al., 2021), which could cause delays for model synchronization and inactivation of some participants, i.e., stragglers, making it challenging to leverage the full potential of FMs in FL settings.
@@
Trustworthiness challenges emphasize the concerns regarding privacy, security, and ethical considerations in the lifecycle of FedFM, from the pre-training and model adaptation to the application stages. We present two representative challenges from this perspective: (1) intellectual property: Intellectual Property (IP) protection in FedFM primarily involves attributing ownership rights for both models and data. From the server’s perspective, broadcasting a pre-trained model to multiple nodes for fine-tuning poses IP protection and security risks (e.g., model theft), necessitating measures to safeguard IP rights and ensure model integrity (Kang et al., 2025); (2) privacy leakage: Although FL does not immediately share data, studies have shown that it may not always guarantee sufficient privacy preservation (Geiping et al., 2020), as model parameters (e.g., weights or gradients) may leak sensitive information to malicious adversaries (Zhu et al., 2019). (3)Poisoning Attacks: FL systems are inherently vulnerable to attacks due to their wide attack surface and reliance on network communication (Li et al., 2024c). Poisoning attacks are carried out by malicious participants, aiming to bias the global model to the desire of attackers.
In the following sections (Sections 4-6), we will review and discuss the existing work with respect to efficiency, adaptability, and trustworthiness.
@@
Figure 1. Taxonomy of research in foundation models with federated learning.
4. Efficiency
The remarkable capabilities of FMs are often accompanied by substantial resource demands, posing significant challenges regarding efficiency of FedFM. To address these bottlenecks, research has increasingly prioritized resource-efficient approaches. In this section, we review and discuss the current state of the art in efficiency, focusing on three primary streams: parameter-efficient fine-tuning, model compression, and zeroth-order optimization. We further break down each category into sub-areas for more detailed analysis.
4.1. Parameter-Efficient Fine-Tuning